TU Next Question 4 of 21 (1 point) Use Table A.2 to find the z-score for which the area to its left is 0.52. The z score for the given area is

media%2F819%2F819f3294-d77b-4a34-9e0a-42

media%2F464%2F4649251c-cf9a-49a9-a449-fb

media%2F0ac%2F0acbd4c2-b551-4484-af73-aa

media%2F45c%2F45c51609-04f6-4bd6-bfbd-07
Answer #1

4)
P(Z<z) 52%
z= NORMSINV(0.52)
z= 0.050153583

5)
P(Z<z1) 14%
z1= NORMSINV(0.14)
z1= -1.080319341

P(Z<z2) 86%
z2= NORMSINV(0.86)
z2= 1.080319341

8)
P(Z<z) 10%
z= NORMSINV(0.1)
z= -1.281551566
I know that, z = (X-mean)/sd
(X-mean)/sd = -1.2816
X= -1.2816*9.5+80.1
X= 67.92

10)
P( X>140)
= 1 - P(X<140)

I know that, z = (X-mean)/(sd)
z1 = (140-110.6)/23)
z1= 1.2783

hence,
P( X>140)
=1- P(Z<1.2783)
1 - NORMSDIST(1.2783)
0.100572

11)
P( X>47)
= 1 - P(X<47)

I know that, z = (X-mean)/(sd)
z1 = (47-39)/4)
z1= 2.0000

hence,
P( X>47)
=1- P(Z<2)
1 - NORMSDIST(2)
0.022750